### v1

@@ 2.4 SCD mappings
Let us denote by ${\cal Z}_{nm}$ the metric space of all n-dimensional subspaces of $\mathbb{R}^{n+m}$ equipped with the metric
$d_{{\cal Z}_{nm}}(L_{1},L_{2}):=\|P_{1}-P_{2}\|,$
where $P_{i}$ is the symmetric $(n+m)\times(n+m)$ matrix representing the orthogonal projection onto $L_{i},i=1,2$.
For notational reasons, we usually identify $\mathbb{R}^{n+m}$ with $\mathbb{R}^{n}\times\mathbb{R}^{m}$, i.e., for a subspace $L\in{\cal Z}_{nm}$ we denote its elements by $(u,v)$ instead of $\begin{pmatrix}u\\ v\end{pmatrix}$. Analogously, given two matrices $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{m\times n}$ so that $L={\rm rge\;}\begin{pmatrix}A\\ B\end{pmatrix}$, we write $L={\rm rge\;}(A,B)$.
Lemma 2.9.
cf.[9, Lemma 3.1] The metric space ${\cal Z}_{nm}$ is (sequentially) compact.
With each $L\in{\cal Z}_{nm}$ one can associate its adjoint subspace $L^{*}$ defined by
@@
Since $\dim L^{\perp}=m$, it follows that $L^{*}\in{\cal Z}_{mn}$ (i.e., its dimension is $m$). It is easy to see that
$L^{*}=S_{nm}L^{\perp},\mbox{ where }S_{nm}=\left(\begin{array}[]{lc}0&-I_{m}\\ I_{n}&0\end{array}\right).$
Further we have $(L^{*})^{*}=L$ and $d_{{\cal Z}_{mn}}(L_{1}^{*},L_{2}^{*})=d_{{\cal Z}_{nm}}(L_{1},L_{2})$ for all subspaces $L_{1},L_{2}\in{\cal Z}_{nm}$ and thus the mapping $L\to L^{*}$ is an isometry between ${\cal Z}_{nm}$ and ${\cal Z}_{mn}$, cf. [9]. In what follows, the symbol $L^{*}$ signifies both the adjoint subspace to some $L\in{\cal Z}_{nm}$ as well as an arbitrary subspace from ${\cal Z}_{mn}$. This double role, however, cannot lead to a confusion.
@@
We say that $F$ is graphically smooth of dimension $n$ at $(\bar{x},\bar{y})$ if $T_{\mathrm{gph}\,F}(\bar{x},\bar{y})\in{\cal Z}_{nm}$. By $\mathcal{O}_{F}$ we denote the subset of $\mathrm{gph}\,F$, where $F$ is graphically smooth of dimension $n$.
@@
Next we introduce the two derivative-like mappings ${\mathcal{S}}F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows{\cal Z}_{nm}$ and ${\mathcal{S}}^{*}F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows{\cal Z}_{mn}$ defined by
${\mathcal{S}}F(x,y):$ =$=\{L\in{\cal Z}_{nm}\,\big{|}\,\exists(x_{k},y_{k}){\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}(x,y)\mbox{ such that }\lim_{k\to\infty}d_{{\cal Z}_{nm}}\big{(}L,T_{\mathrm{gph}\,F}(x_{k},y_{k})\big{)}=0\},$
@@
Note that ${\mathcal{S}}F$ coincide with the mapping $\mathrm{cl}\,\widehat{\mathcal{S}}F$, where
$\widehat{\mathcal{S}}F(x,y):=\begin{cases}T_{\mathrm{gph}\,F}(x,y)&\mbox{if $(x,y)\in{\cal O}_{F}$},\\ \emptyset&\mbox{else.}\end{cases}$
In particular, $\mathrm{gph}\,{\mathcal{S}}F$ is closed in $\mathbb{R}^{n}\times\mathbb{R}^{m}\times{\cal Z}_{nm}$ and, due to the isometry $L\leftrightarrow L^{*}$, $\mathrm{gph}\,{\mathcal{S}}^{*}F$ is closed in $\mathbb{R}^{n}\times\mathbb{R}^{m}\times{\cal Z}_{mn}$ as well.
@@
By virtue of the definition, the SCD property at $(\bar{x},\bar{y})$ is obviously equivalent with the condition ${\mathcal{S}}F(\bar{x},\bar{y})\not=\emptyset$.
Since we consider convergence in the compact metric space ${\cal Z}_{nm}$, by using similar arguments as in the proof of [8, Lemma 3.6], we readily obtain the following result.
Lemma 2.12.
A mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ has the SCD property at $(x,y)\in\mathrm{gph}\,F$ if and only if $(x,y)\in\mathrm{cl}\,{\cal O}_{F}$. Further, $F$ is an SCD mapping if and only if $\mathrm{cl}\,{\cal O}_{F}=\mathrm{cl}\,\mathrm{gph}\,F$, i.e., $F$ is graphically smooth of dimension $n$ at the points of a dense subset of its graph.
For SCD mappings, the adjoint SC derivative ${\mathcal{S}}^{*}F$ forms some kind of skeleton for the limiting coderivative $D^{*}F$. It contains the information which is important in applications, in particular for numerical computations, but is much easier to compute than the limiting coderivative. On the other hand, the limiting coderivative is a more general object important for arbitrary set-valued mappings.
@@
For the sake of completeness we state also the following calculus rules.
Proposition 2.14 (cf.[9, Theorem 4.1]).
Given a mapping $G:\mathbb{R}^{l}\rightrightarrows\mathbb{R}^{m}$ and a mapping $\Phi:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{l}\times\mathbb{R}^{m}$, consider the mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ given by
$\mathrm{gph}\,F=\{(x,y)\,\big{|}\,\Phi(x,y)\in\mathrm{gph}\,G\}.$
Further assume that we are given $(x,y)\in\mathrm{gph}\,F$ such that $\Phi$ is continuously differentiable in some neighborhood of $(x,y)$, $\nabla\Phi(x,y)$ has full row rank and $G$ has the SCD property at $\Phi(x,y)$. Then $F$ has the SCD property at $(x,y)$ and
${\mathcal{S}}F(x,y)=\{L\in{\cal Z}_{nm}\,\big{|}\,\nabla\Phi(x,y)L\in{\mathcal{S}}G(\Phi(x,y))$
${\mathcal{S}}^{*}F(x,y)=S_{nm}\nabla\Phi(x,y)^{T}S_{lm}^{T}{\mathcal{S}}^{*}G(\Phi(x,y)).$
Proposition 2.15.
Let $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ have the SCD property at $(x,y)\in\mathrm{gph}\,F$ and let $h:\Omega\to\mathbb{R}^{m}$ be continuously differentiable at $x\in\Omega$ where $\Omega\subseteq\mathbb{R}^{n}$ is open. Then $F+h$ has the SCD property at $(x,y+h(x))$ and
${\mathcal{S}}(F+h)(x,y+h(x))=\Big{\{}\left(\begin{matrix}I_{n}&0\\ \nabla h(x)&I_{m}\end{matrix}\right)L\,\big{|}\,L\in{\mathcal{S}}F(x,y)\Big{\}},$
${\mathcal{S}}^{*}(F+h)(x,y+h(x))=\Big{\{}\left(\begin{matrix}I_{m}&0\\ \nabla h(x)^{T}&I_{n}\end{matrix}\right)L^{*}\,\big{|}\,L^{*}\in{\mathcal{S}}^{*}F(x,y)\Big{\}}.$
Proof.
Follows from Proposition 2.14 and the observation that $\mathrm{gph}\,(F+h)=\{(x,y)\,\big{|}\,(x,y-h(x))\in\mathrm{gph}\,F\}$. ∎
An important class of SCD mappings is constituted by graphically Lipschitzian mappings:
Definition 2.16 (cf.[28, Definition 9.66]).
@@
From Proposition 2.14 and Lemma 2.13 one can now easily derive the following result:
Lemma 2.17 (cf.[9, Corollary 4.3]).
@@
${\mathcal{S}}F(\bar{x},\bar{y})=\{\nabla\Phi(\bar{x},\bar{y})^{-1}{\rm rge\;}(I,A)\,\big{|}\,A\in\overline{\nabla}f(\bar{u})\},$
${\mathcal{S}}^{*}F(\bar{x},\bar{y})=\{S_{nm}\nabla\Phi(\bar{x},\bar{y})^{T}S^{T}_{nm}{\rm rge\;}(I,A^{T})\,\big{|}\,A\in\overline{\nabla}f(\bar{u})\},$
@@
Prominent examples for graphically Lipschitzian mappings are the subdifferential mapping $\partial f$ for prox-regular and subdifferentially continuous functions $f:\mathbb{R}^{n}\to(-\infty,\infty]$, cf. [28, Proposition 13.46], or more generally, locally maximally hypomonotone mappings, cf. [8].
Theorem 2.18.
Assume that the mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ has the SCD property at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$. If $F$ is strictly proto-differentiable at $(\bar{x},\bar{y})$ then ${\mathcal{S}}F(\bar{x},\bar{y})$ is a singleton. Further, if $F$ is graphically Lipschitzian of dimension $n$ at $(\bar{x},\bar{y})$ then the reverse implication also holds true.
Proof.
Assume that $F$ is strictly proto-differentiable, then we know by Lemma 2.4 that $\bar{L}:=T_{\mathrm{gph}\,F}(\bar{x},\bar{y})$ is a subspace. Further, from (4) we conclude that
Hence, $\bar{n}:=\dim\bar{L}\geq n$ and if $\bar{n}=n$ then ${\mathcal{S}}F(\bar{x},\bar{y})=\{\bar{L}\}$ follows. Now assume that $\bar{n}>n$ and choose an orthogonal basis $(u_{i},v_{i})$, $i=1,\ldots,\bar{n}$, of $\bar{L}$. Since $(\bar{x},\bar{y})\in(\mathrm{cl}\,{\cal O}_{F})\setminus{\cal O}_{F}$, there must be a sequence $(x_{k},y_{k}){\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}(\bar{x},\bar{y})$ satisfying $(x_{k},y_{k})\not=(\bar{x},\bar{y})$ $\forall k$. Taking into account (4), for every $k$ sufficiently large and every $i=1,\ldots,\bar{n}$ we can find elements $(u_{i}^{k},v_{i}^{k})\in L_{k}:=T_{\mathrm{gph}\,F}(x_{k},y_{k})$ with $\|(u_{i}^{k},v_{i}^{k})-(u_{i},v_{i})\|<1/\sqrt{\bar{n}}$. We claim now that for those $k$ the vectors $(u_{i}^{k},v_{i}^{k})$, $i=1,\ldots,\bar{n}$, are linearly independent. Assume on the contrary, that for some large $k$ these vectors are linearly dependent and we can find reals $\lambda_{i}^{k}$, $i=1,\ldots,\bar{n}$ with
$\sum_{i=1}^{\bar{n}}\lambda_{i}^{k}(u_{i}^{k},v_{i}^{k})=0\mbox{ and }\sum_{i=1}^{\bar{n}}(\lambda_{i}^{k})^{2}=1.$
Then by orthogonality of $(u_{i},v_{i})$, $i=1,\ldots,\bar{n}$, together with the Cauchy-Schwarz inequality we obtain that
11 11 $=\|\sum_{i=1}^{\bar{n}}\lambda_{i}^{k}(u_{i},v_{i})\|\leq\|\sum_{i=1}^{\bar{n}}\lambda_{i}^{k}(u_{i}^{k},v_{i}^{k})\|+\|\sum_{i=1}^{\bar{n}}\lambda_{i}^{k}\big{(}(u_{i},v_{i})-(u_{i}^{k},v_{i}^{k})\big{)}\|$
$\leq\Big{(}\sum_{i=1}^{\bar{n}}(\lambda_{i}^{k})^{2}\Big{)}^{1/2}\Big{(}\sum_{i=1}^{\bar{n}}\|(u_{i},v_{i})-(u_{i}^{k},v_{i}^{k})\|^{2}\Big{)}^{1/2}<1\cdot(\sum_{i=1}^{\bar{n}}\frac{1}{\bar{n}})^{1/2}=1,$
a contradiction. Thus $(u_{i}^{k},v_{i}^{k})$, $i=1,\ldots,\bar{n}$, are linearly independent and $\dim L_{k}\geq\bar{n}$ follows. However, by definition of the set ${\cal O}_{F}$ we have $\dim L_{k}=n$, a contradiction. Hence $\bar{n}=n$ and ${\mathcal{S}}F(\bar{x},\bar{y})=\{\bar{L}\}$ holds true.
To show the second assertion, assume that $F$ is graphically Lipschitzian of dimension $n$ at $(\bar{x},\bar{y})$ and let $W$, $\Phi$, $U$ and $f$ be as in Definition 2.16. Note that for all sequences $t_{k}\downarrow 0$, $(x_{k},y_{k}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F\cap W}}(\bar{x},\bar{y})$ and $(\hat{x}_{k},\hat{y}_{k}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F\cap W}}(\bar{x},\bar{y})$ such that
$\lim_{k\to\infty}\frac{(\hat{x}_{k},\hat{y}_{k})-(x_{k},y_{k})}{t_{k}}=:(\xi,\eta)$
exists, we have that $\Phi(x_{k},y_{k})=:(u_{k},f(u_{k})){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,f}}(\bar{u},f(\bar{u}))=\Phi(\bar{x},\bar{y})$, $\Phi(\hat{x}_{k},\hat{y}_{k})=:(\hat{u}_{k},f(\hat{u}_{k})){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,f}}(\bar{u},f(\bar{u}))$ and
$\lim_{k\to\infty}\frac{(\hat{u}_{k},f(u_{k}))-(u_{k},f(u_{k}))}{t_{k}}$ $=\lim_{k\to\infty}\frac{\nabla\Phi(\bar{x},\bar{y})\big{(}(\hat{x}_{k},\hat{y}_{k})-(x_{k},y_{k})\big{)}+o(\|(\hat{x}_{k},\hat{y}_{k})-(x_{k},y_{k})\|)}{t_{k}}$
$=\nabla\Phi(\bar{x},\bar{y})(\xi,\eta).$
@@ 2.5 On semismooth mappings
  1. 1.
    A set $A\subseteq\mathbb{R}^{n}$ is called semismooth at a point $\bar{x}\in A$ if for every $\epsilon>0$ there is some $\delta>0$ such that
    $|\langle x^{*},x-\bar{x}\rangle|\leq\epsilon\|x-\bar{x}\|\|x^{*}\|\ \forall x\in A\cap{\cal B}_{\delta}(\bar{x})\ \forall x^{*}\in N_{A}(x).$
  2. 2.
    A set-valued mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is called semismooth at a point $(\bar{x},\bar{y})\in\mathrm{gph}\,F$, if $\mathrm{gph}\,F$ is semismooth at $(\bar{x},\bar{y})$, i.e., for every $\epsilon>0$ there is some $\delta>0$ such that
    $|\langle x^{*},x-\bar{x}\rangle-\langle y^{*},y-\bar{y}\rangle|$
    $\qquad\leq\epsilon\|(x,y)-(\bar{x},\bar{y})\|\|(y^{*},x^{*})\|\ \forall(x,y)\in\mathrm{gph}\,F\cap{\cal B}_{\delta}(\bar{x},\bar{y})\ \forall(y^{*},x^{*})\in\mathrm{gph}\,D^{*}F(x,y).$ (6)
Note that the semismooth property was defined in [7] in a different but equivalent way. The class of semismooth mappings is rather broad. We mention here two important classes of multifunctions having this property.
Proposition 2.20 (cf.[8, Proposition 2.10]).
  1. (i)
    Every mapping whose graph is the union of finitely many closed convex sets is semismooth at every point of its graph.
  2. (ii)
    Every mapping with closed subanalytic graph is semismooth at every point of its graph.
For single-valued functions there holds the following result.
Proposition 2.21 (cf. [7, Proposition 3.7]).
Assume that $F:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is a single-valued mapping which is Lipschitzian near $\bar{x}$. Then the following two statements are equivalent.
  1. (i)
    $F$ is semismooth at $\bar{x}$.
  2. (ii)
    For every $\epsilon>0$ there is some $\delta>0$ such that
    $\|F(x)-F(\bar{x})-C(x-\bar{x})\|\leq\epsilon\|x-\bar{x}\|\ \forall x\in{\cal B}_{\delta}(\bar{x})\ \forall C\in\mathrm{conv}\,\overline{\nabla}F(x),$
    i.e.,
    $\sup_{C\in\mathrm{conv}\,\overline{\nabla}F(x)}\|F(x)-F(\bar{x})-C(x-\bar{x})\|=o(\|x-\bar{x}\|).$ (7)
Let us compare the semismooth property with the semismooth property as introduced in [26].
Definition 2.22.
Let $\Omega\subseteq\mathbb{R}^{n}$ be nonempty and open. A function $F:\Omega\to\mathbb{R}^{m}$ is semismooth at $\bar{x}\in\Omega$, if it is Lipschitzian near $\bar{x}$ and if
$\lim_{{\textstyle{A\in\mathrm{conv}\,\overline{\nabla}F(\bar{x}+tu^{\prime})\atop u^{\prime}\to u,\ t\downarrow 0}}}Au^{\prime}$
exists for all $u\in\mathbb{R}^{n}$. If F is semismooth at all $\bar{x}\in\Omega$, we call $F$ semismooth on $\Omega$.
Then the following relation is well-known, see, e.g., [31, Proposition 2.7].
Proposition 2.23.
Let $F:\Omega\to\mathbb{R}^{m}$ be defined on the open set $\Omega\subset\mathbb{R}^{n}$. Then for $\bar{x}\in\Omega$ the following two statements are equivalent.
  1. 1.
    $F$ is semismooth at $\bar{x}$,
  2. 2.
    $F$ is Lipschitz near $\bar{x}$, the directional derivative $F^{\prime}(\bar{x};\cdot)$ exists and (7) is fulfilled.
Hence, by virtue of Proposition 2.21, for locally Lipschitz functions the difference between the semismooth property and the semismooth property is just the directional differentiability. Further, we want to mention that under local Lipschitz continuity the semismooth property as defined by Gowda [13] coincides with our semismooth property. However, there are also single-valued functions which are semismooth but not locally Lipschitz, e.g., the function $x\mapsto\sqrt{|x|}$ at 00. This example demonstrates that even for single-valued functions the semismooth property differs from the semismooth property more than only by the lack of directional differentiability.
Sometimes it is useful to replace $\mathrm{gph}\,D^{*}F(x,y)$ in the relation (6) by another set.
@@
Note that the semismooth property introduced in Definition Definition 2.19 amounts to the semismooth property of $F$ at $(\bar{x},\bar{y})$ with respect to $\Gamma(x,y)=\mathrm{gph}\,D^{*}F(x,y)=\{(y^{*},x^{*})\,\big{|}\,(x^{*},-y^{*})\in N_{\mathrm{gph}\,F}(x,y)\}$.
The notion of $\Gamma$-semismoothness has two main applications. Firstly, the calculus rules for coderivatives often yield only inclusions of the form $D^{*}F(x,y)(y^{*})\subset\Sigma(x,y)(y^{*})$. However, as in our motivating application, sometimes it is not necessary to compute the exact limiting coderivative but elements from $\Sigma$ related to the semismooth property. Thus, if we can show that $F$ is $\Gamma$-ss with respect to the mapping $\Gamma(x,y)=\mathrm{gph}\,\Sigma(x,y)$ then we can also work with $\Sigma$ instead of $D^{*}F$.
@@
Proposition 2.26 (cf. [12, Proposition 2.12]).
Consider the mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ and the pair $(\bar{x},\bar{y})\in\mathrm{gph}\,F$. Then the following two statements are equivalent:
@@
    For every $\epsilon>0$ there is some $\delta>0$ such that for every $(x,y)\in\mathrm{gph}\,F\cap{\cal B}_{\delta}(\bar{x},\bar{y})$ and every $L\in{\mathcal{S}}F(x,y)$ there holds
    ${\rm dist}((x-\bar{x},y-\bar{y}),L)\leq\epsilon\|(x-\bar{x},y-\bar{y})\|.$

### v2

@@ 2.4 SCD mappings
Let us denote by ${\cal Z}_{n,m}$ the metric space of all $n$-dimensional subspaces of $\mathbb{R}^{n+m}$ equipped with the metric
$d_{{\cal Z}_{n,m}}(L_{1},L_{2}):=\|P_{1}-P_{2}\|,$
where $P_{i}$ is the symmetric $(n+m)\times(n+m)$ matrix representing the orthogonal projection onto $L_{i},i=1,2$. By [13, Lemma 3.1], ${\cal Z}_{n,m}$ is (sequentially) compact.
For notational reasons, we usually identify $\mathbb{R}^{n+m}$ with $\mathbb{R}^{n}\times\mathbb{R}^{m}$, i.e., for a subspace $L\in{\cal Z}_{n,m}$ we denote its elements by $(u,v)$ instead of $\begin{pmatrix}u\\ v\end{pmatrix}$. Analogously, given two matrices $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{m\times n}$ so that $L={\rm rge\;}\begin{pmatrix}A\\ B\end{pmatrix}$, we write $L={\rm rge\;}(A,B)$.
With each $L\in{\cal Z}_{n,m}$ one can associate its adjoint subspace $L^{*}$ defined by
@@
Since $\dim L^{\perp}=m$, it follows that $L^{*}\in{\cal Z}_{m,n}$ (i.e., its dimension is $m$). Further we have $(L^{*})^{*}=L$ and $d_{{\cal Z}_{m,n}}(L_{1}^{*},L_{2}^{*})=d_{{\cal Z}_{n,m}}(L_{1},L_{2})$ for all subspaces $L_{1},L_{2}\in{\cal Z}_{n,m}$ and thus the mapping $L\to L^{*}$ is an isometry between ${\cal Z}_{n,m}$ and ${\cal Z}_{m,n}$, cf. [13]. In what follows, the symbol $L^{*}$ signifies both the adjoint subspace to some $L\in{\cal Z}_{n,m}$ as well as an arbitrary subspace from ${\cal Z}_{m,n}$. This double role, however, cannot lead to a confusion.
@@
We say that $F$ is graphically smooth of dimension $n$ at $(\bar{x},\bar{y})$ if $T_{\mathrm{gph}\,F}(\bar{x},\bar{y})\in{\cal Z}_{n,m}$. By $\mathcal{O}_{F}$ we denote the subset of $\mathrm{gph}\,F$, where $F$ is graphically smooth of dimension $n$.
@@
Next we introduce the two derivative-like mappings ${\mathcal{S}}F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows{\cal Z}_{n,m}$ and ${\mathcal{S}}^{*}F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows{\cal Z}_{m,n}$ defined by
${\mathcal{S}}F(x,y):$ =$=\{L\in{\cal Z}_{n,m}\,\big{|}\,\exists(x_{k},y_{k}){\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}(x,y)\mbox{ such that }\lim_{k\to\infty}d_{{\cal Z}_{n,m}}\big{(}L,T_{\mathrm{gph}\,F}(x_{k},y_{k})\big{)}=0\},$
@@
Note that ${\mathcal{S}}F$ coincides with the mapping $\mathrm{cl}\,\widehat{\mathcal{S}}F$, where
$\widehat{\mathcal{S}}F(x,y):=T_{\mathrm{gph}\,F}(x,y),\ (x,y)\in\mathrm{dom}\,\widehat{\mathcal{S}}F:={\cal O}_{F}.$
In particular, $\mathrm{gph}\,{\mathcal{S}}F$ is closed in $\mathbb{R}^{n}\times\mathbb{R}^{m}\times{\cal Z}_{n,m}$ and, due to the isometry $L\leftrightarrow L^{*}$, $\mathrm{gph}\,{\mathcal{S}}^{*}F$ is closed in $\mathbb{R}^{n}\times\mathbb{R}^{m}\times{\cal Z}_{m,n}$ as well.
@@
By virtue of the definition, the SCD property at $(\bar{x},\bar{y})$ is obviously equivalent with the condition ${\mathcal{S}}F(\bar{x},\bar{y})\not=\emptyset$. Since we consider convergence in the compact metric space ${\cal Z}_{n,m}$ we readily obtain that $F$ has the SCD property at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$ if and only if $(\bar{x},\bar{y})\in\mathrm{cl}\,{\cal O}_{F}$.
For SCD mappings, the adjoint SC derivative ${\mathcal{S}}^{*}F$ forms some kind of skeleton of the limiting coderivative $D^{*}F$. It contains the information which is important in applications, in particular for numerical computations, but is much easier to compute than the limiting coderivative. On the other hand, the limiting coderivative is a more general object important for arbitrary set-valued mappings.
@@
We refer to [12, 13] for several calculus rules for SCD mappings. An important class of SCD mappings is constituted by graphically Lipschitzian mappings:
Definition 2.10 (cf. [35, Definition 9.66]).
@@
Lemma 2.11 (cf. [13, Corollary 4.3]).
@@
${\mathcal{S}}F(\bar{x},\bar{y}$ )=$)=\{\nabla\Phi(\bar{x},\bar{y})^{-1}{\rm rge\;}(I,A)\,\big{|}\,A\in\overline{\nabla}f(\bar{u})\},$
${\mathcal{S}}^{*}F(\bar{x},\bar{y})$ =$=\{S_{nm}\nabla\Phi(\bar{x},\bar{y})^{T}S^{T}_{nm}{\rm rge\;}(I,A^{T})\,\big{|}\,A\in\overline{\nabla}f(\bar{u})\}\mbox{ with }S_{nm}:=\begin{pmatrix}0&-I_{m}\\ I_{n}&0\end{pmatrix},$
@@
Prominent examples for graphically Lipschitzian mappings are the subdifferential mapping $\partial f$ for prox-regular and subdifferentially continuous functions $f:\mathbb{R}^{n}\to(-\infty,\infty]$, cf. [35, Proposition 13.46] or, more generally, locally maximally hypomonotone mappings, cf. [12].
@@ 2.5 On semismooth mappings
A set-valued mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is called semismooth at a point $(\bar{x},\bar{y})\in\mathrm{gph}\,F$, if for every $\epsilon>0$ there is some $\delta>0$ such that
$|\langle x^{*},x-\bar{x}\rangle-\langle y^{*},y-\bar{y}\rangle|$ (2.8)
$\quad\leq\epsilon\|(x,y)-(\bar{x},\bar{y})\|\|(y^{*},x^{*})\|\ \forall(x,y)\in\mathrm{gph}\,F\cap{\cal B}_{\delta}(\bar{x},\bar{y})\ \forall(y^{*},x^{*})\in\mathrm{gph}\,D^{*}F(x,y).$
Note that the semismooth property was defined in [11] in a different but equivalent way. The class of semismooth mappings is rather broad. For instance, every mapping whose graph is either the union of finitely many convex sets or a closed subanalytic set, is semismooth at every point of its graph [12, Proposition 2.10].
A locally Lipschitz mapping $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, is semismooth at $\bar{x}$ if and only if $F$ is G-semismooth at $\bar{x}$ [11, Proposition 3.7]. However, there are also single-valued mappings which are semismooth but neither locally Lipschitz and nor G-semismooth, e.g., the function $x\mapsto\sqrt{|x|}$ at 00.
Sometimes it is useful to replace $\mathrm{gph}\,D^{*}F(x,y)$ in the relation (2.8) by another set.
@@
Note that the semismooth property introduced in Definition 2.12 amounts to the semismooth property of $F$ at $(\bar{x},\bar{y})$ with respect to $\Gamma(x,y)=\mathrm{gph}\,D^{*}F(x,y)$.
The notion of $\Gamma$-semismoothness has two main applications. Firstly, the calculus rules for coderivatives often yield only inclusions of the form $D^{*}F(x,y)(y^{*})\subset\Sigma(x,y)(y^{*})$ with some closed-graph mapping $\Sigma$. However, as in our motivating application, sometimes it is not necessary to compute the exact limiting coderivative but elements from $\Sigma$ related to the semismooth property. Thus, if we can show that $F$ is $\Gamma$-ss with respect to the mapping $\Gamma(x,y)=\mathrm{gph}\,\Sigma(x,y)$ then we can also work with $\Sigma$ instead of $D^{*}F$.
@@
Proposition 2.15 (cf. [9, Proposition 2.12]).
Assume that $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ has the SCD property around the pair $(\bar{x},\bar{y})\in\mathrm{gph}\,F$. Then the following two statements are equivalent:
@@
    $\limsup\limits_{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}\ \sup\limits_{L\in{\mathcal{S}}F(x,y)}{{\rm dist}((x-\bar{x},y-\bar{y}),L)}/{\|(x-\bar{x},y-\bar{y})\|}=0.$